OpenStudy (anonymous):

how bout 4n^2-9=0 n=

OpenStudy (anonymous):

4n^2=9

OpenStudy (anonymous):

Yes, now to isolate the n you must divide by what?

OpenStudy (eyust707):

This is just like the last one!

OpenStudy (eyust707):

add 9 to both sides then divide both side by 4

OpenStudy (anonymous):

And once you are done with that step, you can remove the square by squaring both sides. Remmeber the plus and minus that goes with the square root.

OpenStudy (anonymous):

so we are looking at n^2/4=9/4

OpenStudy (eyust707):

nope the 4 wouldnt be there anymore on the left side

OpenStudy (anonymous):

ok so 9/4

OpenStudy (eyust707):

n^2 = 9/4

OpenStudy (eyust707):

then sqrt both sides... remember the + or - sign too

OpenStudy (anonymous):

this is differant then the last one since the numerator is more correct

OpenStudy (anonymous):

will it be a mixed number or -9/4,9/4

OpenStudy (eyust707):



OpenStudy (eyust707):

sqrt both sides